# On the Effectiveness of Microservices Tactics and Patterns to Reduce Energy Consumption: An Experimental Study on Trade-Offs

### v1

@@ Abstract
Index Terms:
microservices, tactics, patterns, energy consumption, performance, maintainability, controlled experiment
@@ I Introduction
Over the last few years, microservices have more and more replaced traditional monolithic systems for cloud-based applications. A microservice-based system decomposes its functionality into smaller, independent services that are connected through lightweight communication protocols [1]. Many companies like Amazon and Netflix have adopted microservices in their cloud infrastructures [2]. Typical advantages of microservices are improved maintainability via firm module boundaries, ease of deployment, and scalability [3].
@@
One less frequently discussed quality attribute (QA) for microservices is environmental sustainability, with energy efficiency as the most important sub-characteristics. The growing concern for sustainable software is a global phenomenon, and industry must determine how to best move towards green software and green IT [4]. However, there is currently a substantial gap between microservices-related research on environmental sustainability and industrial needs [5]. Similar to the whole ICT industry, microservices practitioners are pressured to improve environmental sustainability due to three reasons. First, they (will) have to comply with new or incoming legislation regarding sustainability [6]. Second, environmentally unsustainable software often increases business spending on operations and maintenance, such as data center cooling. Many clusters are so power-consuming that they are put in regions with cheap electricity, i.e., it is also economically advantageous to enhance the energy efficiency of microservice-based systems. Lastly, the ICT industry contributes to environmental pollution via greenhouse gas emissions due to its energy demand [7], which both employees and customers can perceive negatively.
@@
Typical techniques to improve QAs are architectural tactics and patterns. Bachmann et al. [8] explain architectural tactics as strategic approaches or methods to design software systems to achieve specific quality goals. This involves making design decisions to directly influence system-level QAs. In our case, the main quality goal is energy consumption. Osses et al. [9] define architectural patterns as abstract descriptions of a system’s structure and the behavior within that structure. Patterns provide a blueprint for organizing the components of a system in a way that addresses particular recurring problems. In practice, it is also essential to be aware of potential trade-offs with other QAs when applying tactics and patterns.
@@
While some techniques to improve the energy efficiency of microservices have started to emerge [5], empirical evidence for their effectiveness is still unclear. Tactics and patterns are often documented without knowing their concrete effects and downsides [10], e.g., without conducting experiments. Sometimes, evidence from controlled environments about the impact of such techniques also does not carry over to industrial contexts [11]. Having reliable empirical evidence for the effectiveness of such energy techniques, as well as potential trade-offs with other QAs, is very important for microservices practitioners trying to improve environmental sustainability.
@@ II Background and Related Work
In this section, we present related work about microservices tactics and patterns, as well as their energy efficiency. Microservices represent an architectural style that emphasizes encapsulating business-specific functionality within fine-granular services that communicate with lightweight mechanisms [12]. The single responsibility of each service enables independent testing, deployment, and scaling [2]. Due to its widespread adoption in industry, microservices architecture has attracted significant attention in academic research.
@@
Two influential concepts when designing microservice-based systems are architectural tactics and patterns, and several literature studies try to provide an overview. Osses et al. [9] conducted a systematic literature review, identifying 44 patterns and tactics for microservices across 8 papers. In their review, they proposed a taxonomy of microservices patterns with five categories. While they discussed the QAs associated with the found patterns, energy consumption or efficiency were not mentioned. Similarly, Valdivia et al. [13] conducted a systematic literature review about the relationship between 34 microservices patterns and 6 QAs, again without energy-related considerations. They reported that clearly defined relationships are often lacking, highlighting the need for more empirical research. In a third systematic literature review, Li et al. [14] focused on microservices tactics and the associated improved QAs. They identified 19 tactics that address 6 QAs in microservices architecture, again neglecting energy consumption. While these reviews summarize the available microservices tactics and patterns, the detailed QA impact of applying these techniques is still an understudied area. Vale et al. [10] tried to alleviate this via an interview study with 9 microservices practitioners. They synthesized a mapping of 14 patterns and their perceived positive and negative QA impact. Quantitative empirical evidence has also started to accumulate in the area of evolvability [15], performance [16], reliability [17], and understandability [18]. Nonetheless, the energy consumption impact of microservices patterns and tactics remains an open question.
@@
However, the energy consumption of microservices has been studied from other perspectives than tactics and patterns, with a recent review by Araújo et al. [5] curating 37 primary studies. While this is not a large number and not all of these papers directly focused on energy consumption, it is reasonable to assume that the environmental sustainability of microservices will become an increasingly prominent quality concern among practitioners. To date, it has still received comparatively limited attention in research and practice. One example is a proactive scheduling model proposed by Fontana de Nardin et al. [19] to reduce energy consumption with minimal impact on performance through resource elasticity heuristics. Their evaluation with a microservice-based application in a private cloud achieved a 27.9% reduction in energy consumption. Similarly, Berry et al. [20] investigated the effect of resource elasticity by comparing horizontal scaling of a monolithic application to its microservices version. The microservices version outperformed the monolithic one in energy use when configured for optimal availability and performance. Lastly, Cortellessa et al. [21] applied a genetic algorithm to generate architectural deployment alternatives of microservices with optimal trade-offs between performance, deployment cost, and power consumption. They found a significant negative impact on response time when a power consumption objective was introduced. Overall, broad empirical evidence on the effectiveness and trade-offs of energy techniques for microservices is unavailable.
@@ III Study Design
To start the accumulation of evidence to fill this gap, we formed an academia-industry collaboration between VU Amsterdam and the Software Improvement Group (SIG) [^1] https://www.softwareimprovementgroup.com, a software consultancy firm specialized in software quality with roughly 160 employees. Together, we first analyzed the microservices literature to identify suitable tactics and patterns to evaluate. Afterward, we conducted a controlled experiment [22]. In this section, we describe the design of this experiment. For transparency and reproducibility, we also published our experiment artifacts on Zenodo. [^2] https://doi.org/10.5281/zenodo.12730144
@@
To keep the scope for the trade-off analysis manageable (RQ2), we focused on two quality attributes that are generally seen as drivers for using a microservice architecture, performance and maintainability [23]. A second reason was that performance and energy consumption are sometimes considered as a trade-off [21].
@@ III-A Experiment Objects
T1 – Distribute pods to different nodes to take advantage of node properties: nodes vary in properties such as energy source, specialized hardware, and workloads from time to time. These differences are crucial when strategically distributing pods across nodes for optimal load or energy balancing [24]. We tested this tactic from the load-balancing perspective, but not from the energy-balancing perspective, as the nodes of our experiment server are based on the same electricity grid. The number of nodes remained unchanged compared to the control group. We used the Kubernetes Scheduler [^3] https://kubernetes.io/docs/concepts/scheduling-eviction/kube-scheduler to dynamically assign pods to nodes based on their resource requirements and dependency relationship.
@@
T2 – Distribute microservices to different containers based on affinity: this energy tactic uses containerized grouping to place several microservices in one container, e.g., services dependent on each other or using the same tech stack [25]. We implemented grouping based on programming languages and resolved library dependencies accordingly.
@@
T3 – Use different proxies for different demands: this tactic proposes to use the most efficient proxy that is suitable for the workload of the system [26], which should reduce resource usage. We tested this tactic with Istio [^4] https://istio.io/latest/about/service-mesh and kube proxy [^5] https://kubernetes.io/docs/reference/command-line-tools-reference/kube-proxy, two of the most popular microservices proxies, which are suitable for different workloads [26]. The experiment baseline used kube proxy and the treatment Istio.
@@
P1 – Backends for Frontends (BFF): this pattern involves creating separate orchestration services tailored to the needs of individual clients or user interfaces [27]. This allows optimizing each backend’s functionality, performance, and development process without impacting other client experiences. It can decrease, e.g., the resources sent to the frontend, the styling description files, and the unused functionality programs on some frontends, making it potentially energy-saving [28]. We implemented two BFFs in the treatment version: one for mobile clients and one for desktop clients. The mobile frontend is more straightforward in terms of interface, which means that fewer resources will be prepared by the corresponding backend. Each frontend received one half of the requests.
@@
P2 – Request Bundle: this pattern describes a batched API request approach to bundle multiple small individual API requests together [29]. It aims to reduce network congestion and hardware resource utilization. The reduced number of requests and responses may also simultaneously reduce the energy consumption [28]. We implemented request bundles for API endpoints with three responses at most, with the system waiting for three requests at a time.
@@
P3 – Caching of Read Requests: caching is a frequently used software performance technique that can also be applied to microservices [30]. If requests often read similar data, caching can eliminate the need to communicate with other services or databases. While it aims to mitigate network congestion and improve API throughput, the Green Software Foundation also lists it as a green software pattern [31]. We implemented a simple in-memory cache per service that stored recent requests, without adding a third-party tool like Redis.
@@ III-B Experimental Materials
To execute the experiment, we used an experiment server cluster in VU Amsterdam’s Green Lab [32, 33], which is specifically designed for energy efficiency experiments. Using this local environment enabled us to collect energy consumption readings and remove unknown infrastructure configurations. The used machines had an Intel Xeon Silver 4208 CPU @ 2.10 GHz and 395 GB of RAM. Intel CPUs offer Running Average Power Limit (RAPL), which is a feature that allows to fetch accurate energy consumption readings from CPUs and DRAM [34]. Many more advanced tools are built on top of RAPL. For our microservices experiment, we chose Scaphandre [^8] https://github.com/hubblo-org/scaphandre, which integrates well with Kubernetes and the Grafana and Prometheus monitoring stack. This also allowed generating intuitive graphs to compare different system versions.
@@ III-C Experiment Variables
The most important independent variable for our experiment was the applied tactic or pattern. In addition to the baseline without any modifications, this included the discussed three tactics (T1-T3) and three patterns (P1-P3), i.e., this variable had a total of seven levels. The second consciously manipulated independent variable was the load we imposed on the system to study different usage scenarios. Load in such systems is typically characterized by user limit and spawn rate. User limit denotes the number of concurrent users accessing the system, while spawn rate denotes the rate at which new users are generated [25]. To introduce decent variety, we decided on three load levels: low, medium, and high. Following Buzato and Goldman [25] in using real website traffic for representative load testing, we extrapolated from the monthly traffic of popular websites reported by Semrush [^12] https://www.semrush.com. For high, we used one-tenth of the number of visits of Amazon [^13] https://www.amazon.com, the most visited retail website in the world, i.e., 175 requests per second (the real number was too large for our experiment cluster). We set the low level to the same number of visits as Bol [^14] https://www.bol.com, a local retail website in the Netherlands, i.e., 52 requests per second. For medium, we then simply used the average of low and high, i.e., 114 request per second. To nicely align these numbers with our experiment trial length of 60 seconds, we rounded the levels to 60, 120, and 180 respectively.
@@ III-D Experiment Design and Execution
Our experiment was based on 7 system versions (1 baseline, 3 tactics, 3 patterns) and 3 load levels (low, medium, high). Using a full factorial design [22], this resulted in 21 experiment configurations. To avoid potential random confounders in the experiment environment, we repeated each configuration 30 times, leading to a total of 630 trials. This was achieved by executing a randomized sequence of the 21 configurations 30 times, which had the additional benefit of minimizing potential infrastructure-level optimizations that might adapt to a fixed sequence. Between each trial, we set a resting time of 1 minute so that the machine could cool down for more reliable energy measurements [35]. Before executing the experiment, we also turned off any unnecessary processes on the machine.
@@ III-E Data Analysis
After all experiment trials were finished, we collected the measurements from the respective tools and started with exploratory data analysis, e.g., by calculating descriptive statistics and creating visualizations like boxplots. We also used the interquartile range method [36] with the standard $1.5$ factor to detect and remove outliers. Afterward, we analyzed if the data we wanted to use for hypothesis testing conformed to a normal distribution via the Shapiro-Wilk test [37].
Since all data of dependent variables turned out to be normally distributed (p-value $>0.05$), we selected the t-test as a suitable method [38]. Based on the described hypotheses, we used a one-tailed test for energy consumption and a two-tailed test for performance. To combat the multiple comparisons problem due to testing many hypotheses [39], we applied the Holm-Bonferroni correction [40] to adjust the corresponding p-value for each test. For identified significant differences (adjusted p-value $<0.05$), we used Cohen’s $d$ [41] to estimate the effect size. According to Cohen, values between $0.0$ and $0.2$ indicate a negligible effect, $0.2$ to $0.5$ a small effect, $0.5$ to $0.8$ a moderate effect, and $0.8$ and above a large effect.

### v2

@@ I Introduction
Over the last few years, microservices have more and more replaced traditional monolithic systems for cloud-based applications. A microservice-based system decomposes its functionality into smaller, independent services that are connected through lightweight communication protocols [alshuqayran2016systematic]. Many companies like Amazon and Netflix have adopted microservices in their cloud infrastructures [thones2015microservices]. Typical advantages of microservices are improved maintainability via firm module boundaries, ease of deployment, and scalability [10220070].
@@
One less frequently discussed quality attribute (QA) for microservices is environmental sustainability, with energy efficiency as the most important sub-characteristics. The growing concern for sustainable software is a global phenomenon, and industry must determine how to best move towards green software and green IT [Verdecchia2021]. However, there is currently a substantial gap between microservices-related research on environmental sustainability and industrial needs [araujo2024energy]. Similar to the whole ICT industry, microservices practitioners are pressured to improve environmental sustainability due to three reasons. First, they (will) have to comply with new or incoming legislation regarding sustainability [GSFLegislation2023]. Second, environmentally unsustainable software often increases business spending on operations and maintenance, such as data center cooling. Many clusters are so power-consuming that they are put in regions with cheap electricity, i.e., it is also economically advantageous to enhance the energy efficiency of microservice-based systems. Lastly, the ICT industry contributes to environmental pollution via greenhouse gas emissions due to its energy demand [Fonseca2019], which both employees and customers can perceive negatively.
@@
Typical techniques to improve QAs are architectural tactics and patterns. bachmann2003deriving explain architectural tactics as strategic approaches or methods to design software systems to achieve specific quality goals. This involves making design decisions to directly influence system-level QAs. In our case, the main quality goal is energy consumption. osses2018exploratory define architectural patterns as abstract descriptions of a system’s structure and the behavior within that structure. Patterns provide a blueprint for organizing the components of a system in a way that addresses particular recurring problems. In practice, it is also essential to be aware of potential trade-offs with other QAs when applying tactics and patterns.
@@
While some techniques to improve the energy efficiency of microservices have started to emerge [araujo2024energy], empirical evidence for their effectiveness is still unclear. Tactics and patterns are often documented without knowing their concrete effects and downsides [Vale2022], e.g., without conducting experiments. Sometimes, evidence from controlled environments about the impact of such techniques also does not carry over to industrial contexts [funke2024experimental]. Having reliable empirical evidence for the effectiveness of such energy techniques, as well as potential trade-offs with other QAs, is very important for microservices practitioners trying to improve environmental sustainability.
@@ II Background and Related Work
In this section, we present related work about microservices tactics and patterns, as well as their energy efficiency. Microservices represent an architectural style that emphasizes encapsulating business-specific functionality within fine-granular services that communicate with lightweight mechanisms [Newman2015]. The single responsibility of each service enables independent testing, deployment, and scaling [thones2015microservices]. Due to its widespread adoption in industry, microservices architecture has attracted significant attention in academic research.
@@
Two influential concepts when designing microservice-based systems are architectural tactics and patterns, and several literature studies try to provide an overview. osses2018exploratory conducted a systematic literature review, identifying 44 patterns and tactics for microservices across 8 papers. In their review, they proposed a taxonomy of microservices patterns with five categories. While they discussed the QAs associated with the found patterns, energy consumption or efficiency were not mentioned. Similarly, Valdivia2019 conducted a systematic literature review about the relationship between 34 microservices patterns and 6 QAs, again without energy-related considerations. They reported that clearly defined relationships are often lacking, highlighting the need for more empirical research. In a third systematic literature review, Li2021 focused on microservices tactics and the associated improved QAs. They identified 19 tactics that address 6 QAs in microservices architecture, again neglecting energy consumption. While these reviews summarize the available microservices tactics and patterns, the detailed QA impact of applying these techniques is still an understudied area. Vale2022 tried to alleviate this via an interview study with 9 microservices practitioners. They synthesized a mapping of 14 patterns and their perceived positive and negative QA impact. Quantitative empirical evidence has also started to accumulate in the area of evolvability [Bogner2019-PeerJ], performance [Pinciroli2023], reliability [ElMalki2023], and understandability [Bogner2024]. Nonetheless, the energy consumption impact of microservices patterns and tactics remains an open question.
@@
However, the energy consumption of microservices has been studied from other perspectives than tactics and patterns, with a recent review by araujo2024energy curating 37 primary studies. While this is not a large number and not all of these papers directly focused on energy consumption, it is reasonable to assume that the environmental sustainability of microservices will become an increasingly prominent quality concern among practitioners. To date, it has still received comparatively limited attention in research and practice. One example is a proactive scheduling model proposed by FontanadeNardin2021 to reduce energy consumption with minimal impact on performance through resource elasticity heuristics. Their evaluation with a microservice-based application in a private cloud achieved a 27.9% reduction in energy consumption. Similarly, Berry2024 investigated the effect of resource elasticity by comparing horizontal scaling of a monolithic application to its microservices version. The microservices version outperformed the monolithic one in energy use when configured for optimal availability and performance. Lastly, Cortellessa2024 applied a genetic algorithm to generate architectural deployment alternatives of microservices with optimal trade-offs between performance, deployment cost, and power consumption. They found a significant negative impact on response time when a power consumption objective was introduced. Overall, broad empirical evidence on the effectiveness and trade-offs of energy techniques for microservices is unavailable.
@@ III Study Design
To start the accumulation of evidence to fill this gap, we formed an academia-industry collaboration between VU Amsterdam and the Software Improvement Group (SIG) [^1] https://www.softwareimprovementgroup.com, a software consultancy firm specialized in software quality with roughly 160 employees. Together, we first analyzed the microservices literature to identify suitable tactics and patterns to evaluate. Afterward, we conducted a controlled experiment [Wohlin2024]. In this section, we describe the design of this experiment. For transparency and reproducibility, we also published our experiment artifacts on Zenodo. [^2] https://doi.org/10.5281/zenodo.12730144
@@
To keep the scope for the trade-off analysis manageable (RQ2), we focused on two quality attributes that are generally seen as drivers for using a microservice architecture, performance and maintainability [Bogner2019]. A second reason was that performance and energy consumption are sometimes considered as a trade-off [Cortellessa2024].
@@ III-A Experiment Objects
T1 – Distribute pods to different nodes to take advantage of node properties: nodes vary in properties such as energy source, specialized hardware, and workloads from time to time. These differences are crucial when strategically distributing pods across nodes for optimal load or energy balancing [valera2022pisco]. We tested this tactic from the load-balancing perspective, but not from the energy-balancing perspective, as the nodes of our experiment server are based on the same electricity grid. The number of nodes remained unchanged compared to the control group. We used the Kubernetes Scheduler [^3] https://kubernetes.io/docs/concepts/scheduling-eviction/kube-scheduler to dynamically assign pods to nodes based on their resource requirements and dependency relationship.
@@
T2 – Distribute microservices to different containers based on affinity: this energy tactic uses containerized grouping to place several microservices in one container, e.g., services dependent on each other or using the same tech stack [ContainerizedGrouping]. We implemented grouping based on programming languages and resolved library dependencies accordingly.
@@
T3 – Use different proxies for different demands: this tactic proposes to use the most efficient proxy that is suitable for the workload of the system [IstioImpact], which should reduce resource usage. We tested this tactic with Istio [^4] https://istio.io/latest/about/service-mesh and kube proxy [^5] https://kubernetes.io/docs/reference/command-line-tools-reference/kube-proxy, two of the most popular microservices proxies, which are suitable for different workloads [IstioImpact]. The experiment baseline used kube proxy and the treatment Istio.
@@
P1 – Backends for Frontends (BFF): this pattern involves creating separate orchestration services tailored to the needs of individual clients or user interfaces [BackendsForFrontends]. This allows optimizing each backend’s functionality, performance, and development process without impacting other client experiences. It can decrease, e.g., the resources sent to the frontend, the styling description files, and the unused functionality programs on some frontends, making it potentially energy-saving [Rani2024]. We implemented two BFFs in the treatment version: one for mobile clients and one for desktop clients. The mobile frontend is more straightforward in terms of interface, which means that fewer resources will be prepared by the corresponding backend. Each frontend received one half of the requests.
@@
P2 – Request Bundle: this pattern describes a batched API request approach to bundle multiple small individual API requests together [RequestBundle]. It aims to reduce network congestion and hardware resource utilization. The reduced number of requests and responses may also simultaneously reduce the energy consumption [Rani2024]. We implemented request bundles for API endpoints with three responses at most, with the system waiting for three requests at a time.
@@
P3 – Caching of Read Requests: caching is a frequently used software performance technique that can also be applied to microservices [Gottlieb2023]. If requests often read similar data, caching can eliminate the need to communicate with other services or databases. While it aims to mitigate network congestion and improve API throughput, the Green Software Foundation also lists it as a green software pattern [GreenSoftwareFoundation_Cache_Static_Data]. We implemented a simple in-memory cache per service that stored recent requests, without adding a third-party tool like Redis.
@@ III-B Experimental Materials
To execute the experiment, we used an experiment server cluster in VU Amsterdam’s Green Lab [procaccianti_green_2015, mendez_ten_2024], which is specifically designed for energy efficiency experiments. Using this local environment enabled us to collect energy consumption readings and remove unknown infrastructure configurations. The used machines had an Intel Xeon Silver 4208 CPU @ 2.10 GHz and 395 GB of RAM. Intel CPUs offer Running Average Power Limit (RAPL), which is a feature that allows to fetch accurate energy consumption readings from CPUs and DRAM [RAPL]. Many more advanced tools are built on top of RAPL. For our microservices experiment, we chose Scaphandre [^8] https://github.com/hubblo-org/scaphandre, which integrates well with Kubernetes and the Grafana and Prometheus monitoring stack. This also allowed generating intuitive graphs to compare different system versions.
@@ III-C Experiment Variables
The most important independent variable for our experiment was the applied tactic or pattern. In addition to the baseline without any modifications, this included the discussed three tactics (T1-T3) and three patterns (P1-P3), i.e., this variable had a total of seven levels. The second consciously manipulated independent variable was the load we imposed on the system to study different usage scenarios. Load in such systems is typically characterized by user limit and spawn rate. User limit denotes the number of concurrent users accessing the system, while spawn rate denotes the rate at which new users are generated [ContainerizedGrouping]. To introduce decent variety, we decided on three load levels: low, medium, and high. Following ContainerizedGrouping in using real website traffic for representative load testing, we extrapolated from the monthly traffic of popular websites reported by Semrush [^12] https://www.semrush.com. For high, we used one-tenth of the number of visits of Amazon [^13] https://www.amazon.com, the most visited retail website in the world, i.e., 175 requests per second (the real number was too large for our experiment cluster). We set the low level to the same number of visits as Bol [^14] https://www.bol.com, a local retail website in the Netherlands, i.e., 52 requests per second. For medium, we then simply used the average of low and high, i.e., 114 request per second. To nicely align these numbers with our experiment trial length of 60 seconds, we rounded the levels to 60, 120, and 180 respectively.
@@ III-D Experiment Design and Execution
Our experiment was based on 7 system versions (1 baseline, 3 tactics, 3 patterns) and 3 load levels (low, medium, high). Using a full factorial design [Wohlin2024], this resulted in 21 experiment configurations. To avoid potential random confounders in the experiment environment, we repeated each configuration 30 times, leading to a total of 630 trials. This was achieved by executing a randomized sequence of the 21 configurations 30 times, which had the additional benefit of minimizing potential infrastructure-level optimizations that might adapt to a fixed sequence. Between each trial, we set a resting time of 1 minute so that the machine could cool down for more reliable energy measurements [cruz2021green]. Before executing the experiment, we also turned off any unnecessary processes on the machine.
@@ III-E Data Analysis
After all experiment trials were finished, we collected the measurements from the respective tools and started with exploratory data analysis, e.g., by calculating descriptive statistics and creating visualizations like boxplots. We also used the interquartile range method [vinutha2018detection] with the standard $1.5$ factor to detect and remove outliers. Afterward, we analyzed if the data we wanted to use for hypothesis testing conformed to a normal distribution via the Shapiro-Wilk test [hanusz2016shapiro].
Since all data of dependent variables turned out to be normally distributed (p-value >0.05>0.05), we selected the t-test as a suitable method [Student1908]. Based on the described hypotheses, we used a one-tailed test for energy consumption and a two-tailed test for performance. To combat the multiple comparisons problem due to testing many hypotheses [barnett2022multiple], we applied the Holm-Bonferroni correction [holm1979simple] to adjust the corresponding p-value for each test. For identified significant differences (adjusted p-value <0.05<0.05), we used Cohen’s $d$ [cohen1988statistical] to estimate the effect size. According to Cohen, values between $0.0$ and $0.2$ indicate a negligible effect, $0.2$ to $0.5$ a small effect, $0.5$ to $0.8$ a moderate effect, and $0.8$ and above a large effect.